OpenStudy (decarr432):

Let f(x)=(9x^3-20)^2 and g(x)=9x^3-20. Given that f(x)=(h^og)(x), find h(x)

OpenStudy (danjs):

f(x) = h(g(x)) (9x^3-20) ^2 = h( 9x^3-20) how does h change that 9x^3 - 20 into the f(x)

OpenStudy (decarr432):

it makes all of f(x) squared?

OpenStudy (mathmale):

Sorry, but your f(x)=(h^og)(x) is incorrectly written. Please fix it.

OpenStudy (decarr432):

How so?

OpenStudy (decarr432):

Whats incorrect?

OpenStudy (decarr432):

All that is given if that \ f(x)=(h^og)(x)

OpenStudy (danjs):

right so the h(x) is just squares the input h(x)=x^2

OpenStudy (mathmale):

f(x)=(h^og)(x) why that "^ " symbol here?

OpenStudy (decarr432):

because I don't know how to put the o between h and g

OpenStudy (decarr432):

thats all it is

OpenStudy (mathmale):

" ^ " does not belong in there. I think you mean f(x) = (h o g)(x).

OpenStudy (danjs):

(h o g)(x), there is a character for it but i forget how to put it in

OpenStudy (mathmale):

f(x) = (h o g)(x) is easily readable and is mathematically correct. This is what you want to calculate. What does " (h o g)(x) " signify to you? What's its name, and what does it ask of you?

OpenStudy (decarr432):

I have no clue bud I'm sorry

OpenStudy (decarr432):

Just started it and it gives no info on it

OpenStudy (mathmale):

I believe I used the word "composition" the last time I helped you. Remember that? What math operations are required by (h o g) ?

OpenStudy (mathmale):

"I begin with function h(x); I throw out "x" from both sides of h(x) and replace "x" with .... what? "

OpenStudy (mathmale):

Can you complete that sentence? I want you to understand what an algebraic composite function is. I went back and looked at our previous conversation and see that I did mention "composition" then.

OpenStudy (decarr432):

I'm lost sorry

OpenStudy (decarr432):

i'm going to close so people don't get spammed

OpenStudy (mathmale):

Sorry, Decarr. You'll need to decide for yourself whether or not you want to learn the meaning of algebraic compositions, and if you do, to make the necessary effort.\

OpenStudy (decarr432):

I am sorry, I'm just slow at picking up on math

OpenStudy (decarr432):

Oh okay I just looked up the definition I think I get it now

OpenStudy (mathmale):

" (h o g) " denotes a composition. Functions h(x) and g(x) are given. To find this composition, you focus on h(x). You throw out the "x" and replace it with g(x). In other words, g(x) becomes the input to function h(x). If you take calculus later, you'll see this "composition of functions" again and again.

OpenStudy (decarr432):

I get what you were trying to tell me I'm sorry I didn't understand but I do now

OpenStudy (mathmale):

In this particular problem, f(x) = (h o g)(x). f(x) and g(x) are given, but h(x) is not. Our job is to determine what the function h(x) looks like.

OpenStudy (mathmale):

Please go to the very top of this conversation. Look at f(x) and g(x). What relationship do you see between them?

OpenStudy (decarr432):

9x^3-20 the only difference is one is squared

OpenStudy (mathmale):

EXACTLY RIGHT. So, based upon this, take a guess. How is h(x) definited?

OpenStudy (decarr432):

h(x)=9x^3-20?

OpenStudy (mathmale):

As you said before, the only difference betweeen f(x) and g(x) is that the latter is squared. Is that right? Is that your meaning?

OpenStudy (decarr432):

Yeah

OpenStudy (mathmale):

Let's try this: find the composition of h(x) = x^2 and g(x)= 9x^3-20.

OpenStudy (mathmale):

That's written as (h o g)(x), same as before.

OpenStudy (mathmale):

Use g(x) as the input to h(x), as before.

OpenStudy (mathmale):

Decarr: we are close to finishing this. OpenStudy could be wrong in describing what you are doing now, but says you are "just looking around." I need your ongoing attention if we're to finish this problem.

OpenStudy (decarr432):

Im here im thinking of how to put it together

OpenStudy (decarr432):

would it be like x(9x^3-20)^2?

OpenStudy (decarr432):

because the example that I looked up shows kinda like that

OpenStudy (mathmale):

type out g(x), please.

OpenStudy (decarr432):

g(x)=x(9x^3-20)^2

OpenStudy (mathmale):

But at the very top of this conversation, g(x) is definited as 9x^3 - 20. h(x) is x^2. I ask you to replace the "x" in h(x) with g(x) and the "x" in x^2 with (9x^3-20). By doing this you are forming the composition (h o g) (x).

OpenStudy (mathmale):

h(x) is definited as h(x) = x^2. throw out the x on both sides. Obtain h( ) = ( )^2. What goes inside those parentheses?

OpenStudy (mathmale):

Again, you are forming the composition (h o g)(x). It will be ( )^2 after you write 9x^3 - 20 inside those parentheses.

OpenStudy (decarr432):

So simple? (9x^3-20)^2 or no x?

OpenStudy (decarr432):

And thanks for putting up with me I know im frustrating

OpenStudy (mathmale):

NO on the left. Right. Now compare your result with f(x). Are your result and f(x) the same or different?

OpenStudy (mathmale):

I meant NO "x" on the left. Correct.

OpenStudy (mathmale):

(h o g)(x) = (9x^3 - 20)^2. that's all. Is this the same as f(x) or not?

OpenStudy (mathmale):

Decarr: May I have your attention, please?

OpenStudy (decarr432):

Sorry had to make baby sister a bottle for bed

OpenStudy (decarr432):

And yes it is the same

OpenStudy (mathmale):

Please let me know if you have to leave our conversation. Otherwise I sit here waiting for you without knowing the reason.

OpenStudy (decarr432):

I apologize

OpenStudy (mathmale):

All right. That goes to show that h(x) = x^2 is correct. We're done.

OpenStudy (mathmale):

The algebraic operation we used again and again here is called " ? "

OpenStudy (decarr432):

Composition right?

OpenStudy (mathmale):

yes. very good. Hope you have a better idea of what "composition" means in algebra, and can correctly explain expressions such as (h o g)(x).

OpenStudy (decarr432):

Yes thanks again happy new year

OpenStudy (mathmale):

Same to you. Take care, see you again.